OpenStudy (josee):

how close to -3 do we have to take x for the inequality 1/(x+3)^2 >700 -within at least 0.097 -within at least 0.057 -within at least 0.037 -within at least 0.077 -within at least 0.017

OpenStudy (amistre64):

1/(x+3)^2 >700-.097

OpenStudy (amistre64):

1/(700 - (error)) = (x+3)^2 x = \(-3\pm\sqrt{\cfrac{1}{700-(error)}}\)

OpenStudy (josee):

its not .097

OpenStudy (amistre64):

hmm .... we might have to just tack it on the end then :)

OpenStudy (amistre64):

\[-3\pm\sqrt{\cfrac{1}{700}}\pm(error)\] maybe?

OpenStudy (josee):

what do you put in for error?

OpenStudy (amistre64):

its just a guess, but id say the given requirements: -within at least 0.097 -within at least 0.057 -within at least 0.037 -within at least 0.077 -within at least 0.017